### v1

# From Demonstrations to Safe Deployment: Path-Consistent Safety Filtering for Diffusion Policies

Abstract
[…]
I Introduction
One of the main goals of robotics is to reduce the mental and physical burden on human workers by automating tedious and repetitive tasks. Recently, imitation learning using diffusion models [ho2020denoising] and flow matching [lipman2022flow] has enabled robots to successfully perform these complex, long-horizon manipulation tasks [chi2023diffusion, reuss2024multimodal]. These advances have been reinforced by the availability of large-scale demonstration datasets [o2024open] and the emergence of vision-language-action models [black2024pi_0, shukor2025smolvla]. These policies, which we collectively refer to as diffusion policies (DPs) in this work, can perform tasks in dynamic environments [ye2025ra, wang2025one]. However, since DPs are black-box models, they lack the required safety guarantees for avoiding collisions or safely interacting with dynamic objects. This prevents their deployment in human-centric environments, where safety has to be formally guaranteed.
Figure 1: Deploying DPs in dynamic environments with moving objects requires safeguarding mechanisms, as the intended policy actions may be unsafe. Reactive strategies, such as control barrier functions, often drive the agent into out-of-distribution (OOD) states not seen during training, leading to unpredictable behavior. We propose that safety mechanisms for DPs should remain consistent with the robot’s intended path to avoid out-of-distribution states and preserve high task success rates.
Many works have shown how safety can be ensured for learning-based policies during deployment [brunke2022safe, krasowski2023provably], for example, using control barrier functions [ames_2019_ControlBarrier, singletary2022safety] or predictive safety filters [wabersich_2021_PredictiveSafety, tearle2021predictive]. Most of these methods are reactive safety mechanisms that satisfy safety constraints by changing the path of the robot away from the dynamic object. However, the vast amount of available demonstrations [o2024open] is not recorded with any of such safety mechanisms in place. Therefore, as illustrated in Figure 1, reactive safety mechanisms often drive the system away from the demonstration data distribution into unseen states. DPs are particularly prone to errors in these out-of-distribution cases, increasing the risk of task failure [agia2025unpacking, xu2025can, romer2025failure]. To avoid this out-of-distribution situation, the interventions of a safety mechanism for DPs should ideally keep the system close to the desired actions of the policy.
We achieve this by proposing a Path-consistent safety filter (PACS) for DPs. The core idea is to use path-consistent braking to guarantee safety while keeping the system on its intended path. For this purpose, PACS translates the entire action sequence (chunk) generated by the policy into an intended trajectory, along which the robot is slowed down or stopped. We use set-based reachability analysis [althoff_2019_Effortless] to verify the safety constraints, allowing for real-time deployment at $1\text{\,}\mathrm{kHz}$.
In summary, our main contributions are:
  • We perform the first provably safe deployment of DPs for human-robot interaction (HRI) tasks, representing safety-critical and challenging examples of dynamic real-world environments.
[…]
II Related Work
Safety Filters in Robotics
Deploying autonomous agents in safety-critical applications requires mechanisms to prevent harmful behavior. A common approach is to apply safety filters, which have been extensively reviewed in recent surveys [brunke2022safe, krasowski2023provably, wabersich_2023_DataDrivenSafety, hsu_2024_SafetyFilter]. Safety filters typically operate by verifying whether a failsafe action exists that can bring the robot to a safe state after executing the action proposed by the policy [hsu_2024_SafetyFilter]. Several techniques follow this approach: Model predictive safety filters [gros_2020_SafeReinforcement, wabersich_2021_PredictiveSafety] and control barrier functions [ames_2019_ControlBarrier, marvi_2021_SafeReinforcement, singletary2022safety, morton_2025_SafeTaskconsistent] formulate an optimization problem to compute the closest safe control input to the desired input. In contrast, methods based on reachability analysis [althoff_2015_IntroductionCORA, akametalu_2014_ReachabilitybasedSafe, fisac_2019_GeneralSafety, krasowski_2020_SafeReinforcement, shao_2021_ReachabilitybasedTrajectory, thumm_2022_ProvablySafe, thumm_2025_GeneralSafety] check whether the system remains outside unsafe states during the execution of the action and failsafe trajectory. These failsafe trajectories can be path-consistent, stopping the robot along its intended path [shao_2021_ReachabilitybasedTrajectory, thumm_2022_ProvablySafe], or non-path-consistent, guiding the robot away from unsafe regions.
Safety of Diffusion Policies
Compared to more classical programming-by-demonstration approaches [zeestraten2016online], the stochastic and end-to-end nature of diffusion policies makes them particularly prone to generating unsafe actions and behaving erratically, especially in unseen situations [agia2025unpacking, xu2025can, romer2025failure]. One way to address this problem is by adding an additional safety layer to the learned policy, for example, using control barrier functions [brunke2025semantically]. Another example is RAIL [jung2024railreachabilityaidedimitationlearning], which maintains a failsafe plan based on reachability analysis, defined as a stopping motion that stays as close as possible to the nominal trajectory. While theoretically well-grounded, such post-processing methods only consider static constraints so far and significantly distort the action distribution, leading to a notable drop in task performance [jung2024railreachabilityaidedimitationlearning]. To address this problem, several works directly incorporate safety constraints into the iterative action generation (denoising) process of diffusion- and flow-based policies. Adding cost gradients [carvalho2025motion, saha2024edmp] or classifier-free guidance [nikken2024denoising] can steer the generated action sequence towards safe regions, but these methods do not guarantee hard constraint satisfaction. As an alternative, iterative projections to a safe set [romer2024safe, romer2025diffusion, liang2025simultaneous] or control barrier functions [xiao2023safediffuser, dai2025safeflowmatchingrobot] have been injected into the denoising process to provide certain safety guarantees. However, these approaches still change the policy, and their high computational costs limit their applicability to low-dimensional systems [romer2024safe, romer2025diffusion] or offline planning [liang2025simultaneous, xiao2023safediffuser, dai2025safeflowmatchingrobot]. In summary, real-time safeguarding of DPs in dynamic environments without compromising task success remains an open problem.
III Problem Statement
We aim to ensure a safe deployment of DPs when operating in close proximity to or interacting with dynamic objects, including humans. More specifically, we consider a robotic system with state $\bm{x}(t)\in\mathcal{X}$ and input $\bm{u}(t)\in\mathcal{U}$ at time $t$ that should perform a given task while adhering to a set of safety constraints. We assume that the state $\bm{x}$ of the system contains the pose of dynamic objects up to a bounded measurement error and that the objects have a bounded input set, e.g., a maximal velocity. We assume the availability of a set of trajectories demonstrated by experts $\mathcal{D}=\{\bm{o}_{k},\bm{a}_{k}\}_{k=0}^{K}$} in the form of observation-action pairs. The observations $\bm{o}\in\mathcal{O}$ can contain camera images, language commands, robot joint angles, and other environment information. The action $\bm{a}\in\mathcal{A}$ is a high-level goal that the robot should reach, e.g., a desired joint configuration. From the expert demonstrations, a DP $\pi:\mathcal{A}^{H}\times\mathcal{O}\rightarrow\mathbb{R}_{+}$ is trained that returns a distribution over action chunks $\bm{A}\in\mathcal{A}^{H}$ of length $H$. In this work, we aim to design a controller $\bm{k}:\mathcal{X}\times\mathcal{A}^{H}\times\mathbb{R}\rightarrow\mathcal{U}$} that takes as input the latest action chunk $\bm{A}$ generated by the policy and returns a safe control input $\bm{u}_{\text{safe}}(t)$.
IV Preliminaries
Figure 2: System overview of our proposed path-consistent safety filter (PACS). The policy, conditioned on visual observations and proprioceptive inputs, generates action chunks that are transformed into a sequence of desired waypoints. From these waypoints, we compute a kinematically and dynamically feasible intended trajectory. PACS continuously monitors this trajectory and applies high-frequency safety filtering using reachability analysis to enforce task-specific safety constraints (e.g., collision avoidance or impact force limits).
This section describes the technical background on DPs and reachability-based safety verification.
IV-A Imitation Learning with Diffusion Policies
The demonstration data $\mathcal{D}$ encodes the desired behavior for solving the task. A DP $\pi$ is trained to approximate the conditional action distribution in the demonstrations [chi2023diffusion]. Once trained, diffusion- and flow-based policies generate new actions by iteratively transforming a random noise sample into the learned target distribution [ho2020denoising, lipman2022flow]. Most DPs return a chunk of $H$ consecutive actions $\bm{A}_{k}=(\bm{a}_{k},\bm{a}_{k+1},\dots,\bm{a}_{k+H-1})$} instead of only the next action to improve the temporal consistency and execution speed [chi2023diffusion, zhao2023learning]. During deployment, we can iteratively sample new action chunks $\bm{A}_{k}\sim\pi(\cdot\,|\,\bm{o}_{k})$}, apply the first $h\leq H$ actions $\bm{a}_{k:k+h-1}$ to the system for a time $\Delta t$, respectively, and replan at timestep k+h{k+h}, i.e., after a time $h\Delta t$ has passed. We refer to $H$ as the action chunk length and $h$ as the action execution steps. This general formulation of imitation learning policies encompasses state-of-the-art generative single-task policies [chi2023diffusion, reuss2024multimodal] and generalist vision-language-action models [black2024pi_0, shukor2025smolvla].
IV-B Safety Constraints for Dynamic Objects
The robot and the objects are considered as dynamic systems, and we define the trajectory of a system at time $t$ for a given initial state $\bm{x}_{0}\in\mathcal{X}_{0}$} and a possible input trajectory $\bm{u}(\cdot)$} by $\chi(t;\bm{x}_{0};\bm{u}(\cdot))$}. Following [althoff_2019_Effortless, p. 5], we formally introduce the reachable set $\mathcal{R}(t)$} of a system at time $t$ for a compact set of possible inputs $\mathcal{U}$ by
$\mathcal{R}(t)\coloneq\{\chi(t;\bm{x}_{0};\bm{u}(\cdot))\,|\,\bm{x}_{0}\in\mathcal{X}_{0},\forall t:\bm{u}(t)\in\mathcal{U}\},$ (1)
and denote all states that are reachable during a time interval as $\mathcal{R}([t_{a},t_{b}])=\bigcup_{t\in[t_{a},t_{b}]}\mathcal{R}(t)$}. We denote the reachable occupancy by $\mathcal{C}(t)$. For the reachable occupancy computation of the robot $\mathcal{C}_{\text{robot}}(t)$, we assume a bounded tracking error of the desired trajectory. This assumption could be replaced by applying more advanced methods as presented in [stefan2025ControllerSynthesis].
Following prior work [althoff_2019_Effortless, liu_2021_OnlineVerification, beckert_2017_multiple_safety], we define a collision between an object and the robot if their reachable occupancies $\mathcal{C}_{\text{object}}(t)$ and $\mathcal{C}_{\text{robot}}(t)$ intersect, i.e., a collision event is defined as
$c_{\text{coll}}(t)\coloneq\mathcal{C}_{\text{object}}(t)\cap\mathcal{C}_{\text{robot}}(t)\not$$\emptyset$. (2)
Depending on the task and object, collisions should either be avoided altogether or restricted to non-harmful, low-force interactions. Accordingly, we define two types of safety constraints in accordance with ISO/TS 15066:2016. Power and force limiting (PFL) targets human-robot collaboration with close interaction and allows contacts with humans as long as the kinetic energy of the robot is below pain and injury thresholds $T_{\text{safe}}(t)$. So, the PFL safety constraint is
$c_{\text{safe},\text{PFL}}(t)$ $\coloneqq\overline{c_{\text{coll}}(t)}\,\lor\big(T_{\text{robot}}(t)\leq T_{\text{safe}}(t)\big).$ (3)
The safe energy threshold $T_{\text{safe}}(t)$ can depend on the collision type (constrained collisions for clamping or unconstrained collisions for free movement), the shape of the robot part in contact (e.g., sharp or blunt edges), or the human body part affected in the worst case [kirschner_2024_constrained, kirschner_2024_unconstrained, thumm_2025_GeneralSafety]. In contrast, speed and separation monitoring (SSM) targets coexistence without close interaction, as it does not permit any contact with humans. If a collision can occur, the robot must be at a complete stop with zero kinetic energy, i.e., $T_{\text{safe}}(t)=0$ in (3).
V Path-Consistent Safety Filter for Diffusion Policies
We now detail PACS, which ensures the safe deployment of pretrained DPs. An overview is provided in Figure 2.
V-A Safety-Permitting Diffusion Policy
To maintain high task performance despite path-consistent safety interventions, we need to design the policy in such a way that slowing down or stopping the robot does not in itself represent an out-of-distribution situation. To achieve this, we only include RGB images $\bm{I}(t)$ and joint angles $\bm{q}(t)$ into the observation $\bm{o}(t)$} and no explicit velocity information, which is a common approach in recent works [black2024pi_0, shukor2025smolvla]. Our approach is applicable to various action spaces, but we restrict our presentation to joint velocity control for clarity. In this case, the actions can be the delta joint positions $\Delta\bm{\hat{q}}_{i}=(\bm{\hat{q}}_{i+1}-\bm{\hat{q}}_{i})$, and an action chunk is given by (Algorithm 1, line 5)
$\bm{A}(t)=(\Delta\bm{\hat{q}}_{0},\dots,\Delta\bm{\hat{q}}_{H-1})\sim\pi(\cdot\,|\,\bm{o}(t)).$ (4)
Here, we use $\hat{\cdot}$ to distinguish the desired actions generated by the policy from the actual safe commands sent to the robot. By integrating the desired joint positions forward in time, starting from the current state $\bm{\hat{q}}_{0}$, we compute a sequence of waypoints
$\bm{\hat{p}}(t)=(\bm{\hat{q}}_{0},\dots,\bm{\hat{q}}_{h}),\quad\text{where}\quad\bm{\hat{q}}_{0}=\bm{q}(t),$ (5)
describing the desired path of the policy (Algorithm 1, line 6).
Algorithm 1 Path-Consistent Safety Filter (PACS) for diffusion policies (DPs) and vision-language-action models (VLAs).
1:DP or VLA $\pi(\cdot)$, chunk length $H$, safety filter horizon $h$, policy timestep $\Delta t$, safety step size $\alpha_{s}$, initial configuration $\bm{q}(0)\in\mathcal{X}_{\mathrm{Safe}}$
[…]
4:while $\neg\textit{success}$ do $\triangleright$ High-level (diffusion) steps
5:  $\bm{o}(t)\leftarrow\textsc{GetObservation}(t)$
6:  $\bm{A}(t)=(\Delta\bm{\hat{q}}_{0},\ldots,\Delta\bm{\hat{q}}_{H-1})\sim\pi(\cdot\,|\,\bm{o}(t))$
7:  $(\bm{\hat{q}}_{0}=\bm{q}(t),\dots,\bm{\hat{q}}_{h})\leftarrow\textsc{Integrate}(\bm{A}(t))$
8:  $\chi_{I}([t,t_{h}])\leftarrow\textsc{ComputeTrajectory}((\bm{\hat{q}}_{0},\ldots,\bm{\hat{q}}_{h}))$
9:  $\textit{shield}\leftarrow\textsc{SetTrajectory}(\chi_{I}([t,t_{h}]))$
10:  for $k\leftarrow 0$ to $\frac{h\Delta t}{\alpha_{s}}$ do $\triangleright$ Low-level safety steps
11:   $\bm{m}_{\mathrm{o}}(t)\leftarrow\textsc{GetObjectMeasurements}(t)$
12:   $\bm{u}_{\mathrm{safe}}(t+\alpha_{s})\leftarrow\textsc{ShieldStep}(t,\,\bm{m}_{\mathrm{o}}(t))$
[…]
14:   $t\leftarrow t+\alpha_{s}$
15:  end for
16:  $\textit{success}\leftarrow\textsc{CheckSuccess}(\bm{o}(t))$
17:  if success then
18:   break
19:  end if
20:end while
V-B Path-Consistent Safety for Diffusion Policies
To ensure that our safety filter does not steer the policy into out-of-distribution states, we adopt a path-consistent safety formulation. The key idea is to guarantee safety by modifying the joint speed $\dot{\bm{q}}(t)$ and acceleration $\ddot{\bm{q}}(t)$ while staying on the intended trajectory. Existing path-consistent safety filters for learning-based policies using reachability analysis [thumm_2025_GeneralSafety, thumm_2022_ProvablySafe] can only take as action input a single goal position with zero velocity. However, DPs and vision-language-action models predict a chunk (4) of multiple consecutive actions together and operate at a relatively low update frequency $1/(h\Delta t)$. Therefore, directly applying single-action approaches [thumm_2025_GeneralSafety, thumm_2022_ProvablySafe] to these policies either leads to a very slow execution by stopping at each waypoint or leaves out intermediate action steps, sacrificing fine-grained control and causing a deviation from the desired behavior.
To construct smooth trajectories that align with the predicted action chunk, we propose an intended trajectory planning module (cf. Figure 2). For each action chunk (4), we compute a time-optimized intended trajectory $\chi_{I}([t,t_{h}])=\big\{\bm{x}(\tau)\,\big|\,\tau\in[t,t_{h}]\}$} (Algorithm 1, line 7), where $\bm{x}(t)=\big(\bm{q}(t),\dot{\bm{q}}(t),\ddot{\bm{q}}(t),\dddot{\bm{q}}(t)\big)$ is the state of the system, by solving the optimization problem
$\min_{t_{h},\,\chi_{I}([t,t_{h}])}\quad$ $t_{h}-t$ (6a)
s.t. $\forall\tau\in[t,t_{h}]:\;\bm{x}(\tau)\in\mathcal{X},$ (6b)
$\forall k\in\{1,\dots,h-1\}:$
$\qquad\exists\,t_{k}\in(t_{k-1},t_{k+1}):~\bm{q}(t_{k})=\bm{\hat{q}}_{k},$ (6c)
$\bm{q}(t_{0}=t)=\bm{\hat{q}}_{0},\qquad\bm{q}(t_{h})=\bm{\hat{q}}_{h},$ (6d)
where the set $\mathcal{X}$ defines robot-specific kinematic and dynamic constraints $\bm{q}_{\min}\leq\bm{q}(t)\leq\bm{q}_{\max}$, $\|\dot{\bm{q}}(t)\|\leq\dot{q}_{\max}$, $\|\ddot{\bm{q}}(t)\|\leq\ddot{q}_{\max}$, and $\|\dddot{\bm{q}}(t)\|\leq\dddot{q}_{\max}$. Adjusting the velocity, acceleration, and jerk limits allows us to adapt the execution speed of the action chunk, which gives us a more fine-grained control over the execution speed of the policy compared to modifying the actions directly. Importantly, our intermediate planning module provides a trajectory (6) describing the desired behavior of the policy, which we forward to our path-consistent safety filter (Algorithm 1, line 8).
The safety filter (Algorithm 1, line 9-14) operates at a high frequency and formally guarantees constraint satisfaction while staying on the desired path (5). Based on the current state of the robot $\bm{x}_{0}=\bm{x}(t)$ and the measured object positions $\bm{m}_{\text{o}}(t)$ (Algorithm 1, line 10), we compute a path-consistent monitored trajectory [thumm_2025_GeneralSafety] with intended step $t_{0},t_{\text{I}}]$ and failsafe trajectory (stopping motion) $t_{\text{I}},t_{\text{F}}]$
$\chi_{\text{M}}(t)=\begin{cases}\chi_{I}\big(t;\,\bm{x}_{0},\bm{u}_{I}(\cdot)\big),&t\in[t_{0},t_{\text{I}}],\\[3.99994pt] \chi_{F}\big(t;\,\bm{x}_{I},\bm{u}_{F}(\cdot)\big),&t\in[t_{\text{I}},t_{\text{F}}].\end{cases}$ (7)
At each low-level safety timestep of duration $\alpha_{\text{s}}$, we verify the monitored trajectory (7) for the task-specific safety constraints as described in Section IV-B. If the verification is successful, we execute the intended trajectory $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{I}}(t)$}; otherwise, we proceed one step on the last successfully verified failsafe trajectory, i.e., $\bm{u}_{\text{safe}}(t)=\bm{u}_{\text{F}}(t)$}. This process is repeated until we receive a new action chunk (4) from the policy after $h\Delta t/\alpha_{\text{s}}$ safety steps. Since this safeguarding mechanism is provably safe by induction, as shown in prior work [thumm_2025_GeneralSafety], PACS enables the safe deployment of DPs.
(a) Sorting: Put the red blocks into the box while a person is taking the green blocks.
(b) Handover: Grab Lego blocks from a hand and put them into the box.
(c) Feeding: Carefully put a fork with food in the mouth and pull the fork back out.
Figure 3: Visualization of our three real-world tasks, which require safe and reactive motion in close proximity to the human body. Sorting represents a coexistence task where no collisions are permitted, whereas Handover and Feeding are collaboration tasks requiring non-harmful, low-force interactions.
VI Evaluation
We conduct extensive experiments in simulation and on hardware to evaluate the real-time safeguarding capabilities of our method. Our code will be released upon acceptance to preserve anonymity. Our quantitative evaluation primarily aims to validate the following hypotheses:
  • H1
    Through behavior cloning from demonstrations alone, DPs cannot learn a safe behavior for interaction with moving objects.
  • H2
    With our path-consistent safeguarding method, task success rate remains close to the unsafe policy ($\pm$ $5\text{\,}\%$) in our conducted experiments.
  • H3
    Computing intended trajectories from entire action chunks leads to higher task success rates than single-action approaches.
  • H4
    PACS is real-time capable and does not reduce task execution speed, unless required to ensure safety.
VI-A Simulation Experiments
TABLE I: Robomimic task success rate results on 100100 rollouts. PACS achieves higher success rates than other safety filters.
Method Safe? Lift Can Square Average
Operational space contr. 1.00 0.99 0.74 0.91
OFF 0.92 0.83 0.34 0.70
Contr. barrier func. [singletary2022safety] 0.11 0.00 0.00 0.04
Single-action - SSM 0.97 0.26 0.00 0.41
Single-action - PFL 0.94 0.33 0.04 0.44
PACS (ours) - SSM 0.97 0.80 0.30 0.69
PACS (ours) - PFL 0.93 0.85 0.38 0.72
First, we evaluate the performance of PACS against an unsafe operational space controller and a control barrier function baseline [singletary2022safety] on the three robomimic [mandlekar2021matters] benchmark tasks Lift, Can, and Square. We integrate these tasks in human-robot-gym [thumm_2024_human_robot_gym] and add a spherical dynamic object that moves in pre-defined patterns on the surface of the table. The goal is to fulfill the original task while ensuring the safety constraints $c_{\text{safe},\text{SSM}}$ (for operational space controller, control barrier function, and SSM) or $c_{\text{safe},\text{PFL}}$ (for PFL). To evaluate H3, we further test a single-action version of PACS that sequentially computes new intended trajectories from each action in the action chunk.
We train the DP with a temporal UNet architecture [chi2023diffusion] on the provided dataset of $200$ proficient human teleoperated demonstrations per task. The policy receives the end effector and object poses of the last two timesteps as observations and predicts desired changes in the end effector pose and gripper position as actions. The operational space controller directly controls the robot in Cartesian space to follow these goals. For all other approaches, we use an inverse kinematics solver [coumans_2016_PyBulletPython] to translate the desired changes in end-effector pose into desired joint state actions and control the robot in joint space. Additional implementation details can be found in Table III. For the operational space controller baseline, we use $h=8$ action execution steps and apply each action for $\Delta t=$50\text{\,}\mathrm{ms}$\$. To evaluate the impact of the safety mechanism and the inverse kinematics translation separately, we also test the performance of our safety shield in OFF mode. Here, we compute the same intended trajectory in joint space, but never execute a failsafe trajectory.
[…]
TABLE II: Constraint formulations for our real-world tasks.
Task Constraint Type Body Part Condition $T_{\text{safe}}$ $\mathrm{J}$]
Sorting SSM Hand - 0.000
Handover PFL Hand Constrained 0.014
Unconstrained 0.265
Feeding PFL Head (Eye) Unconstrained 0.001
VI-B Real-World Setup
Second, we conduct hardware experiments with a Franka FR3 manipulator and consider the example application of HRI, as safety is particularly critical in this area.
VI-B1 Tasks
We consider three tasks that require the robot to move safely in close proximity to humans, as visualized in Figure 3:
  • Sorting: We place two red and three green blocks randomly on the table. The robot must put the red blocks into a box while the human is taking the green blocks. To ensure safety, no collisions between the robot and the human are allowed ($c_{\text{safe},\text{SSM}}$).
  • Handover: The robot must pick a block from the hand of a human partner. To ensure safety, the energy of the impact between the gripper and the hand must not exceed a given energy limit ($T_{\text{safe},\text{hand}}$).
  • Feeding: The robot must put a fork with food into the mouth of the user. To ensure safety, the energy of the impact between the head and the fork must not exceed a given limit ($T_{\text{safe},\text{head}}$). Note that $T_{\text{safe},\text{head}}\ll T_{\text{safe},\text{hand}}$. We use a printed AI-generated face with a cut-out mouth for our quantitative experiments and conduct qualitative tests with a real human.
VI-B2 Safety Constraints
We use the SaRA framework [sven_2022_SaRA] for computing the reachable occupancies and intersection checking in real time. We formally identify the safe energy thresholds in (3) for the different tasks [thumm_2025_GeneralSafety]. Further details and numerical values are provided in Table II. We leverage Ruckig [Ruckig] to compute the intended trajectory (6), using the standard joint limits of the FR3, velocity limits between $1.25\text{\,}$ and $2\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-1}$, an acceleration limit of $10\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-2}$ and a jerk limit of $400\text{\,}\mathrm{rad}\text{\,}{\mathrm{s}}^{-3}$ for each joint.
TABLE III: Implementation details for the DP and vision-language-action model in our experiments.
Hyperparameter DP (sim) DP (real) SmolVLA
Demonstrat. per task 200 50 50
Generative model Diffusion Diffusion Flow Matching
Architecture U-Net U-Net VLM + Transformer
Training steps 64k 20k 40k
Learning rate $\text{1}\times\text{10}^{-\text{4}}$} $\text{1}\times\text{10}^{-\text{4}}$} $\text{1}\times\text{10}^{-\text{4}}$}
Batch size 100 64 64
Action chunk len. $H$ 16 16 50
Action exec. steps $h$ 2 6 6
Inference timesteps 100 (DDPM) 10 (DDIM) 10 (Euler)
Sampling time $\Delta t$ 100 $\mathrm{ms}$ 33 $\mathrm{ms}$ 33 $\mathrm{ms}$
VI-B3 Policies
We deploy two generative policies: DP [chi2023diffusion] that uses denoising diffusion [ho2020denoising] and SmolVLA [shukor2025smolvla] using flow matching [lipman2022flow]. The observations consist of joint angles and the images of a wrist and a third-person camera, and the actions are desired changes in joint angles and the gripper width. We provide details on the training data and the policy implementation in Table III. We use CRISP [pro2025crispcompliantros2] for data recording and for translating the joint commands to motor torques. The SmolVLA policy struggled with Sorting and Feeding, likely due to the high precision required; therefore, we only evaluated it on the Handover task.
VI-B4 Metrics
We report the success rate of completing the task (Success) and the success rate of completing the task while remaining safe at all times (Safe Success) based on $30$ test rollouts. Moreover, we report the proportion of timesteps for which the task-specific safety constraints are violated (Safety Viol.) and the duration of successful rollouts.
VI-C Real-World Results
We deploy the policies with our safety filter OFF and active, similar to the simulation experiments. As shown by the results in Table IV, the policies achieve very similar task success rates of about $80\text{\,}\%$ on average with and without our safety filter, supporting H2. The safeguarded policies do not exhibit constraint violations in our experiments. In contrast, the nominal policies are in an unsafe state $56\text{\,}\%$ of the time with each rollout containing constraint violations, resulting in a Safe Success of $0\text{\,}\%$, which confirms H1. These results highlight the necessity of online safety mechanisms when deploying DPs in dynamic environments and the effectiveness of PACS for this purpose.
TABLE IV: Impact of safeguarding in our real-world experiments. PACS guarantees safe deployment and maintains high task success, resulting in a high Safe Success rate.
Task Policy Safety Filter Success Safe Succ. Safety Viol.
Sorting DP OFF 0.77 0.00 0.67 $\pm$ 0.12
PACS (ours) 0.80 0.80 0.00 $\pm$ 0.00
Handover DP OFF 1.00 0.00 0.32 $\pm$ 0.07
PACS (ours) 0.97 0.97 0.00 $\pm$ 0.00
SmolVLA OFF 0.77 0.00 0.41 $\pm$ 0.15
PACS (ours) 0.80 0.80 0.00 $\pm$ 0.00
[…]
PACS (ours) 0.63 0.63 0.00 $\pm$ 0.00
[…]
PACS (ours) 0.80 0.80 $\bm{\pm} 0.00$
We also compare our path-consistent safety approach with a popular reactive method; control barrier functions [singletary2022safety]. Since control barrier functions are better suited for coexistence than for collaboration tasks, we conduct this comparison for the Sorting task. As can be seen from Table V, PACS achieves a $37\text{\,}\%$ higher success rate than the control barrier function-based filter. To further examine this inferior performance, we visualize the training data distribution and the end effector positions of randomly selected rollouts in Figure 4. While our method keeps the policy in distribution, the control barrier function often pushes the robot into out-of-distribution states from which the policy struggles to recover, leading to a higher failure rate. Considering our simulation results, we find that PACS overall achieves a $51\text{\,}\%$ higher task success rate than the control barrier function safety filter. Besides, we observe that even if the policy manages to recover, this usually takes a few seconds, leading to a higher task completion time on average. The average computation time of one safety step with PACS is $0.20\text{\,}\mathrm{ms}$ compared to $0.64\text{\,}\mathrm{ms}$ for the control barrier function, and recalculating the intended trajectory takes about $5\text{\,}\mathrm{ms}$. This supports H4 and shows that our method can be deployed in real time in highly dynamic environments.
Refer to caption OOD Failures
Figure 4: End effector paths for the Sorting task. The color intensity of the trajectories indicates the velocity, and the intensity of the shaded grey areas visualizes the training distribution. Our safety filter slows down the policy without leaving the desired path when the human is nearby. In contrast, the control barrier function (CBF) pushes the robot away from unsafe states, which often leads to out-of-distribution (OOD) states from which the policy cannot recover.
[…]
TABLE V: Baseline comparison and speed analysis for the Sorting task. PACS achieves higher task performance than control barrier functions and can even perform the task faster than the nominal policy.
Human Safety Filter Success Duration $\mathrm{s}$\$]
w/ human Control barrier function 0.43 35.8 $\pm$ 18.8
PACS (ours) 0.80 32.6 $\pm$ 11.1
w/o human - 0.93 25.2 $\pm$ 3.3
PACS (ours) 0.93 21.7 $\pm$ 3.9
Lastly, we test our safety filter for the Feeding task with a real human. As shown in our accompanying video, the robot slows down when approaching the human, and the safeguarded policy inserts and withdraws the fork carefully and precisely. These experiments highlight the potential of our approach to harness the expressiveness of DPs for highly safety-critical applications, such as healthcare.
VII Conclusion
We propose PACS, a general framework for safely deploying action chunking-based imitation learning policies, such as DPs and vision-language-action models, in dynamic environments. Our insight is that safety interventions should be path-consistent to avoid out-of-distribution situations and maintain high task performance. We achieve this by computing an intended trajectory from the action chunk and monitoring it using reachability analysis. Our experiments show that PACS can guarantee safety in real time and achieve high task success even on challenging real-world HRI tasks, whereas reactive safety filtering with control barrier functions frequently leads the policy into unrecoverable states. Since we do not perform constraint-aware online replanning, which would require conditioning the policy on a constraint parameterization, our method is specifically designed to satisfy dynamic constraints, such as moving obstacles. We consider this work an important step towards deploying generative policies in safety-critical, human-centered environments.